OpenStudy (anonymous):

Solve this system of equations using substitution. Put your answer in ordered pair form, (x, y). x + 2y = 7 2x + y = 8

OpenStudy (anonymous):

x = 7 - 2y 2(7 - 2y) + y = 8 14 - 4y + y = 8 14 - 3y = 8 -3y = -6 y = 2

OpenStudy (anonymous):

x=7-2y therefore 14-4y+y=8 therefore y =2 therfore x=3 (3,2)